# Fully Heteroscedastic Count Regression with Deep Double Poisson Networks

### v4

@@ 3 Deep Double Poisson Networks (DDPN)
where $c$ is a normalizing constant. Let $Z$ denote a random variable with a Double Poisson distribution function, then we say $Z\sim\text{DP}(\mu,\gamma)$. In line with previous work (Zou et al., 2013; Aragon et al., 2018), we assume the moment approximations proposed by Efron (1986): $\mathbb{E}[Z]=\mu$ and $\text{Var}[Z]=\frac{\mu}{\gamma}$. We specify a model,  [^1] For both $\hat{\mu}$ and $\hat{\gamma}$ we apply the log “link” function to ensure positivity and numerical stability. We simply exponentiate whenever $\hat{\mu}_{i}$ or $\hat{\gamma}_{i}$ are needed (i.e., to evaluate the P.M.F.) $[\text{log}\ \hat{\mu}_{i},\text{log}\ \hat{\gamma}_{i}]^{T}=\mathbf{f}_{\mathbf{\Theta}}(\mathbf{x}_{i})$, as follows: let $\mathbf{z}_{i}=\mathbf{f}_{\mathbf{\Theta}_{1:L-1}}(\mathbf{x}_{i})$, be the $d$-dimensional hidden representation of the input $\mathbf{x}_{i}$ produced by the first $L-1$ layers of a neural network. We apply two separate affine transformations to this representation to obtain our distribution parameters: $\text{log}(\hat{\mu}_{i})=\boldsymbol{w}_{\mu}^{T}\mathbf{z}_{i}+b_{\mu}$ and $\text{log}(\hat{\gamma}_{i})=\boldsymbol{w}_{\gamma}^{T}\mathbf{z}_{i}+b_{\gamma}$.
@@ 3.2 Loss Attenuation Dynamics of DDPN
The loss function of a regressor $\mathscr{L}(\hat{\mu}_{i},\hat{\phi}_{i})$, taking as input the predicted mean $\hat{\mu}_{i}\in\mathcal{U}$ and dispersion $\hat{\phi}_{i}\in(0,\infty)$, exhibits learnable attenuation if it admits the form $\mathscr{L}(\hat{\mu}_{i},\hat{\phi}_{i})=d(\hat{\phi}_{i})+a(\hat{\phi}_{i})r(\hat{\mu}_{i},y_{i})$, where $\lim_{\hat{\phi}\to\infty}d(\hat{\phi})=\infty$, $\lim_{\hat{\phi}\to\infty}a(\hat{\phi})=0$, and $r\geq 0$ with equality holding iff $\hat{\mu}_{i}=y_{i}$. We call $d$ the dispersion penalty, $a$ the attenuation factor, and $r$ the residual penalty.
@@ 3.3 $\beta$-DDPN: Controllable Loss Attentuation
Notice that if the predicted inverse dispersion, $\hat{\gamma}_{i}$, is sufficiently small (corresponding to large variance), it can completely zero out $\frac{\partial\mathscr{L}_{i}}{\partial\hat{\mu}_{i}}$ regardless of the current value of $\hat{\mu}_{i}$ Thus, during training, a neural network can converge to (and get “stuck” in) suboptimal solutions wherein poor mean fit is explained away via large uncertainty values. To remedy this behavior, we propose a modified loss function, $\mathscr{L}_{i}^{(\beta)}$, also called the Double Poisson $\beta$-NLL:
@@
[…]
@@ 4.3 Real-World Datasets
Aleatoric Only Poisson DNN 0.664 (0.01) 0.553 (0.01) 1.099 (0.02) 0.851 (0.01) 1.023 (0.04) 0.706 (0.01) 0.818 (0.01) 0.559 (0.00)
@@
Aleatoric + Epistemic (DEs) Poisson DNN 0.650 0.547 1.046 0.817 0.996 0.683 0.823 0.556
@@ A.1 Derivation of the DDPN Objective
If we model our regression targets as $y_{i}\sim\text{DP}(\mu_{i},\gamma_{i})$, where $[\mu_{i},\gamma_{i}]^{T}=\mathbf{f}_{\mathbf{\Theta}}(\mathbf{x}_{i})$, then the Double Poisson NLL for an individual prediction is obtained as follows:
@@ B.2.2 Continuous Ranked Probability Score
[…]
@@ B.3.1 Simulation Experiment (Introductory Figure)
To produce Figure 1, we generate a dataset with the following procedure: First, we sample $x$ from a uniform distribution, $x\sim\texttt{Uniform}(0,2\pi)$. Next, we draw an initial proposal for $y$ from a conflation (Hill, 2011) of five identical Poissons, each with rate parameterized by $\lambda(x)=10\sin(x)+10$. We scale $y$ by $-1$ and shift it by $+30$ to achieve high dispersion at low counts and under-dispersion at high counts while maintaining nonnegativity. We generate 800 points from this process for the training split, as well as 100 for validation and 100 for evaluation.
@@ B.3.5 Amazon Reviews
Amazon Reviews is publicly available at https://cseweb.ucsd.edu/~jmcauley/datasets/amazon_v2/. The “Patio, Lawn, and Garden” subset we employ in this work is hosted at https://datarepo.eng.ucsd.edu/mcauley_group/data/amazon_v2/categoryFilesSmall/Patio_Lawn_and_Garden.csv. Our subset contains 793,966 total reviews, which we divide into a 70/10/20 train/val/test split.
@@ B.4.1 Sharpness
Aleatoric Only Poisson DNN 0.262 (0.00) 0.437 (0.01) 0.263 (0.01) 0.211 (0.00)
@@
Aleatoric + Epistemic Poisson DNN 0.261 0.413 0.253 0.210
@@ Appendix C Proofs and Derivations
    If $y_{i}>0$, we can make the substitution $u=\frac{\hat{\mu}_{i}}{y_{i}}$ and rewrite the residual penalty: $r=uy_{i}-y_{i}-y_{i}\log u=y_{i}(u-1-\log u)$. Since $y_{i}>0$, this yields $r=0\implies(u-1-\log u)=0$. We now claim that $h(u)=(u-1-\log u)=0\implies u=1$. To see this, note that $h^{\prime}(u)=1-\frac{1}{u}$ is negative when $u<1$ and positive when $u>1$. Since $h(u)=0$ when $u=1$, strictly increasing to the right of $u$, and strictly decreasing to the left of $u$, 00 is uniquely achieved by $h$ at $u=1$. So $h=0\implies u=1$. Thus, $u-1-\log u=0\implies u=1\implies\frac{\hat{\mu}_{i}}{y_{i}}=1\implies\hat{\mu}_{i}=y_{i}$.

### v5

@@ 3 Deep Double Poisson Networks (DDPN)
where $c$ is a normalizing constant. Let $Z$ denote a random variable with a Double Poisson distribution function, then we say $Z\sim\text{DP}(\mu,\gamma)$. In line with previous work (Zou et al., 2013; Aragon et al., 2018), we assume the moment approximations proposed by Efron (1986): $\mathbb{E}[Z]=\mu$ and $\text{Var}[Z]=\frac{\mu}{\gamma}$. We specify a model,  [^1] For both $\hat{\mu}$ and $\hat{\gamma}$ we apply the log “link” function to ensure positivity and numerical stability. We simply exponentiate whenever $\hat{\mu}_{i}$ or $\hat{\gamma}_{i}$ are needed (i.e., to evaluate the P.M.F.) $\text{log}\ \hat{\mu}_{i},\text{log}\ \hat{\gamma}_{i}]^{T}=\mathbf{f}_{\mathbf{\Theta}}(\mathbf{x}_{i})$, as follows: let $\mathbf{z}_{i}=\mathbf{f}_{\mathbf{\Theta}_{1:L-1}}(\mathbf{x}_{i})$, be the $d$-dimensional hidden representation of the input $\mathbf{x}_{i}$ produced by the first L1L-1 layers of a neural network. We apply two separate affine transformations to this representation to obtain our distribution parameters: $\text{log}(\hat{\mu}_{i})=\boldsymbol{w}_{\mu}^{T}\mathbf{z}_{i}+b_{\mu}$ and $\text{log}(\hat{\gamma}_{i})=\boldsymbol{w}_{\gamma}^{T}\mathbf{z}_{i}+b_{\gamma}$.
@@ 3.2 Loss Attenuation Dynamics of DDPN
[…]
@@ 3.3 $\beta$-DDPN: Controllable Loss Attentuation
Notice that if the predicted inverse dispersion, $\hat{\gamma}_{i}$, is sufficiently small (corresponding to large variance), it can completely zero out $\frac{\partial\mathscr{L}_{i}}{\partial\hat{\mu}_{i}}$ regardless of the current value of $\hat{\mu}_{i}$ . Thus, during training, a neural network can converge to (and get “stuck” in) suboptimal solutions wherein poor mean fit is explained away via large uncertainty values. To remedy this behavior, we propose a modified loss function, $\mathscr{L}_{i}^{(\beta)}$, also called the Double Poisson $\beta$-NLL:
@@
where $\lfloor\cdot\rfloor$ denotes the stop-gradient operation and we once again average across all prediction / target tuples $(\hat{\mu}_{i},\hat{\gamma}_{i},y_{i})$ in the dataset. With this modification we can effectively temper the loss attenuation behavior of DDPN. We now have partial derivatives $\frac{\partial\mathscr{L}_{i}^{(\beta)}}{\partial\hat{\mu}_{i}}=\left(\hat{\gamma}_{i}^{1-\beta}\right)\left(1-\frac{y_{i}}{\hat{\mu}_{i}}\right)$ and $\frac{\partial\mathscr{L}_{i}^{(\beta)}}{\partial\hat{\gamma}_{i}}=-\frac{1}{2\hat{\gamma}_{i}^{1+\beta}}+\hat{\gamma}_{i}^{-\beta}\left(\hat{\mu}_{i}-y_{i}(1+\log\hat{\mu}_{i}-\log y_{i})\right)$. The Double Poisson $\beta$-NLL is parameterized by $\beta\in[0,1]$, where $\beta=0$ recovers the original Double Poisson NLL and $\beta=1$ corresponds to fitting the mean, $\mu$, with no respect to $\gamma$ (while still performing normal weight updates to fit the value of $\gamma$). Thus, we can consider the value of $\beta$ as providing a smooth interpolation between the natural DDPN likelihood and a more mean-focused loss. For an empirical demonstration of the impact of $\beta$ on DDPN, see Figure 6 in Section 4.5.
@@ 4.3 Real-World Datasets
Aleatoric Only
Poisson DNN 0.664 (0.01) 0.553 (0.01) 1.099 (0.02) 0.851 (0.01) 1.023 (0.04) 0.706 (0.01) 0.818 (0.01) 0.559 (0.00)
@@
Aleatoric + Epistemic (DEs)
Poisson DNN 0.650 0.547 1.046 0.817 0.996 0.683 0.823 0.556
@@ A.1 Derivation of the DDPN Objective
If we model our regression targets as $y_{i}\sim\text{DP}(\mu_{i},\gamma_{i})$, where $\mu_{i},\gamma_{i}]^{T}=\mathbf{f}_{\mathbf{\Theta}}(\mathbf{x}_{i})$, then the Double Poisson NLL for an individual prediction is obtained as follows:
@@ B.2.2 Continuous Ranked Probability Score
$\text{CRPS}(\hat{F}_{i},y_{i})$ $=\int_{-\infty}^{\infty}[\hat{F}_{i}(z)-F_{i}(z)]^{2}\,dz$
$=\int_{-\infty}^{\infty}[\hat{F}_{i}(z)-\mathds{1}_{\{z\geq y_{i}\}}]^{2}\,dz$
$=\int_{-\infty}^{y_{i}}\hat{F}_{i}(z)^{2}\,dz+\int_{y_{i}}^{\infty}[\hat{F}_{i}(z)-1]^{2}\,dz$
@@ B.3.1 Simulation Experiment (Introductory Figure)
To produce Figure 1, we generate a dataset with the following procedure: First, we sample $x$ from a uniform distribution, $x\sim\texttt{Uniform}(0,2\pi)$. Next, we draw an initial proposal for $y$ from a conflation (Hill, 2011) of five identical Poissons, each with rate parameterized by $\lambda(x)=10\sin(x)+10$. We scale $y$ by 1-1 and shift it by +30+30 to achieve high dispersion at low counts and under-dispersion at high counts while maintaining nonnegativity. We generate 800 points from this process for the training split, as well as 100 for validation and 100 for evaluation.
@@ B.3.5 Amazon Reviews
Amazon Reviews is publicly available at https://cseweb.ucsd.edu/˜jmcauley/datasets/amazon_v2/. The “Patio, Lawn, and Garden” subset we employ in this work is hosted at https://datarepo.eng.ucsd.edu/mcauley_group/data/amazon_v2/categoryFilesSmall/Patio_Lawn_and_Garden.csv. Our subset contains 793,966 total reviews, which we divide into a 70/10/20 train/val/test split.
@@ B.4.1 Sharpness
Aleatoric Only
Poisson DNN 0.262 (0.00) 0.437 (0.01) 0.263 (0.01) 0.211 (0.00)
@@
Aleatoric + Epistemic
Poisson DNN 0.261 0.413 0.253 0.210
@@ Appendix C Proofs and Derivations
    If $y_{i}>0$, we can make the substitution $u=\frac{\hat{\mu}_{i}}{y_{i}}$ and rewrite the residual penalty: $r=uy_{i}-y_{i}-y_{i}\log u=y_{i}(u-1-\log u)$. Since $y_{i}>0$, this yields $r=0\implies(u-1-\log u)=0$. We now claim that $h(u)=(u-1-\log u)=0\implies u=1$. To see this, note that $h^{\prime}(u)=1-\frac{1}{u}$ is negative when $u<1$ and positive when $u>1$. Since $h(u)=0$ when $u=1$, strictly increasing to the right of $u$, and strictly decreasing to the left of $u$, 0 is uniquely achieved by $h$ at $u=1$. So $h=0\implies u=1$. Thus, $u-1-\log u=0\implies u=1\implies\frac{\hat{\mu}_{i}}{y_{i}}=1\implies\hat{\mu}_{i}=y_{i}$.